solve for v.
-6(1-5v)=54

Answer by DrBeeee(684)

Given:
(1) (-6)*(1-5*v) = 54, notice that I put () around -6 to help you with the negative number.
Firstly divide both sides of the equation by -6 and get
(2) (-6)*(1-5*v)/(-6) = 54/(-6) or
(3) (1-5*v) = -9, where the (-6)'s on the left side cancelled.
Now remove the () and obtain
(4) 1 - 5*v = -9
Now simplify by adding 9 to each side and get
(5) 1 + 9 - 5*v = -9 +9 or
(6) 10 - 5*v = 0
Now add 5*v to each side and get
(7) 10 - 5*v + 5*v = 0 + 5*v or
(8) 10 = 5*v or
(9) v = 2
Is this correct?
Put v = 2 into (1)
Is ((-6)*(1-5*2) = 54)?
Is ((-6)*(-9) = 54)?
Is (54 = 54)? Yes
Answer: The solution is v = 2.